Do P14 % (Round to two decimal places.) % (Round to two decimal places) 43 05% for 2013.
media%2F156%2F1568652f-e483-452c-be4d-50
media%2Fb6b%2Fb6b51f3a-5bf5-4e7b-ba28-78
Answer #1

{Image was not very clear, I have shown the calculations with the best numbers, though review - especially asset and equity numbers}

Net profit Margin = Net Income/Sales

Return on Assets = Net Income/Total Assets

Return on Equity = Net Income/Total Equity

a) Profit Margin for 2014 = 1,753,496/14,146,886 = 12.39%

b) Profit Margin for 2013 = 1,628,487/13,567,335 = 12.00%

c) Return on Assets for 2014 = 1,753,496/14,679,576 = 11.95%

d) Return on Assets for 2013 = 1,628,487/14,115,127 = 11.54%

e) Return on Equity for 2014 = 1,753,496/4,072,723 = 43.05%

f) Return on Equity for 2013 = 1,628,487/4,378,350 = 37.19%

Of the given options for correct statement - Statement 4 is correct. ROE for 2014 is 43.05%. And since ROE for 2014 > 2013, the performance in 2014 is stronger.